How many kilograms of water must be added to 80kg of a 60% iodine solution in order to dilute a 40% iodine solution?
Answer by stanbon(75887)

How many kilograms of water must be added to 80kg of a 60% iodine solution in order to get a 40% iodine solution?
--------
Equation:
water + water = water
-----
0.40*80 + 1*x = 0.60(80+x)
----
40*80 + 100x = 60*80+ 60x
40x = 20*80
x = 40 kg (amt. of water to add)
====================================
Cheers,
Stan H.